# Oh! We Freeze: Improving Quantized Knowledge Distillation via Signal Propagation Analysis for Large Language Models

### v1

@@ Abstract
Large generative models, such as large language models (LLMs) and diffusion models have as revolutionized the fields of NLP and computer vision respectively. However, their slow inference, high computation and memory requirement makes it challenging to deploy them on edge devices. In this study, we propose a light-weight quantization aware fine tuning technique using knowledge distillation (KD-QAT) to improve the performance of 4-bit weight quantized LLMs using commonly available datasets to realize a popular language use case, on device chat applications. To improve this paradigm of finetuning, as main contributions, we provide insights into stability of KD-QAT by empirically studying the gradient propagation during training to better understand the vulnerabilities of KD-QAT based approaches to low-bit quantization errors. Based on our insights, we propose ov-freeze, a simple technique to stabilize the KD-QAT process. Finally, we experiment with the popular 7B LLaMAv2-Chat model at 4-bit quantization level and demonstrate that ov-freeze results in near float-point precision performance, i.e., less than 0.7% loss of accuracy on Commonsense Reasoning benchmarks.
@@ 1 Introduction
The increasing popularity of large generative neural networks, LLaMA Touvron et al. (2023), OPT Zhang et al. (2022) and diffusion models Dhariwal & Nichol (2021); Ho et al. (2020), has revolutionized the field of machine learning exhibiting exceptional capabilities in generating realistic human-like text and images. However, due to large compute and memory requirements of these models, deploying these models on resource-constrained devices is challenging. Among various compression methods available, quantization Nagel et al. (2021); Krishnamoorthi (2018); Jacob et al. (2018) has proven to be a promising technique to optimize various vision and language models for deployment on resource-constrained devices.
@@ 2 Related Work
Many methods have been proposed in the model efficiency space to solve the problem of neural network quantization. Quantization not only reduces model size but also leverages efficient fixed-point representation over floating-point representations. Most quantization techniques can be categorized either as post-training quantization (PTQ) Nagel et al. (2020; 2019); Dong et al. (2019) or quantization-aware-training (QAT)Bhalgat et al. (2020); Esser et al. (2019); Nagel et al. (2022) methods methods. Although, PTQ are go-to compression techniques Frantar et al. (2023); Yao et al. (2022) for most generative models, these methods suffer huge degradation in performance post quantization for low bitwidth quantization. QAT based approaches are difficult for these models due to lack of access to high quality data and training recipes required for finetuning these models.
In this work, we propose to use knowledge distillation to finetune weight only quantized language models to achieve near floating point performance. LLM-QAT Liu et al. (2023) takes a similar approach but uses teacher generated dataset for finetuning, which is not always feasible due to sensitivity of generated dataset on sampling strategies and compute and memory intensive nature of the generation process. In this work, we use publicly available datasets for the distillation process, and perform ablations to understand vulnerabilities in the multi-head self-attention modules of the LLaMAv2-Chat model.
@@ 3.1 Preliminaries
For our experiments, we use AIMET  [^1] AIMET is a product of Qualcomm Innovation Center, Inc., available on GitHub at https://github.com/quic/aimet Siddegowda et al. (2022) to quantize the models to desired bit-widths and use per-channel symmetric quantization for weights and asymmetric quantization for activations respectively. We set the quantization range of each weight quantizer using a mean squared error (MSE) based criteria minimizing the local MSE between FP and quantized tensors.
@@ 3.2 KD-QAT Setup
To demonstrate the effectiveness of our KD-QAT methodology on an established chat application, we experiment with the popular LLaMAv2-Chat model Touvron et al. (2023), and quantize it to W4A16222WxAy quantization indicates quantizing all the weights and output activations to x and y bits respectively.. For knowledge distillation based QAT, we use full-precision chat model as the teacher and quantized model as our student. Contrary to LLM-QATLiu et al. (2023), we observe that using a weighted sum of cross-entropy and KL divergence for training leads to more stable and better convergence.

### v2

@@ Abstract
Large generative models such as large language models (LLMs) and diffusion models have revolutionized the fields of NLP and computer vision respectively. However, their slow inference, high computation and memory requirement makes it challenging to deploy them on edge devices. In this study, we propose a light-weight quantization aware fine tuning technique using knowledge distillation (KD-QAT) to improve the performance of 4-bit weight quantized LLMs using commonly available datasets to realize a popular language use case, on device chat applications. To improve this paradigm of finetuning, as main contributions, we provide insights into stability of KD-QAT by empirically studying the gradient propagation during training to better understand the vulnerabilities of KD-QAT based approaches to low-bit quantization errors. Based on our insights, we propose ov-freeze, a simple technique to stabilize the KD-QAT process. Finally, we experiment with the popular 7B LLaMAv2-Chat model at 4-bit quantization level and demonstrate that ov-freeze results in near floating point precision performance, i.e., less than 0.7% loss of accuracy on Commonsense Reasoning benchmarks.
@@ 1 Introduction
The increasing popularity of large generative neural networks, LLaMA (Touvron et al. (2023), OPT Zhang et al. (2022)) and diffusion models (Dhariwal & Nichol (2021); Ho et al. (2020)), has revolutionized the field of machine learning exhibiting exceptional capabilities in generating realistic human-like text and images. However, due to large compute and memory requirements of these models, deploying these models on resource-constrained devices is challenging. Among various compression methods available, quantization (Nagel et al. (2021); Krishnamoorthi (2018); Jacob et al. (2018)) has proven to be a promising technique to optimize various vision and language models for deployment on resource-constrained devices.
@@ 2 Related Work
Many methods have been proposed in the model efficiency space to solve the problem of neural network quantization. Quantization not only reduces model size but also leverages efficient fixed-point representation over floating-point representations. Most quantization techniques can be categorized either as post-training quantization (PTQ) (Nagel et al. (2020; 2019); Dong et al. (2019)) or quantization-aware-training (QAT)(Bhalgat et al. (2020); Esser et al. (2019); Nagel et al. (2022)) methods methods. Although, PTQ are go-to compression techniques (Frantar et al. (2023); Yao et al. (2022)) for most generative models, these methods suffer huge degradation in performance post quantization for low bitwidth quantization. QAT based approaches are difficult for these models due to lack of access to high quality data and training recipes required for finetuning these models.
In this work, we propose to use knowledge distillation to finetune weight only quantized language models to achieve near floating point performance. LLM-QAT (Liu et al. (2023)) takes a similar approach but uses teacher generated dataset for finetuning, which is not always feasible due to sensitivity of generated dataset on sampling strategies and compute and memory intensive nature of the generation process. In this work, we use publicly available datasets for the distillation process, and perform ablations to understand vulnerabilities in the multi-head self-attention modules of the LLaMAv2-Chat model.
@@ 3.1 Preliminaries
For our experiments, we use AIMET  [^1] AIMET is a product of Qualcomm Innovation Center, Inc., available on GitHub at https://github.com/quic/aimet (Siddegowda et al. (2022)) to quantize the models to desired bit-widths and use per-channel symmetric quantization for weights and asymmetric quantization for activations respectively. We set the quantization range of each weight quantizer using a mean squared error (MSE) based criteria minimizing the local MSE between FP and quantized tensors.
@@ 3.2 KD-QAT Setup
To demonstrate the effectiveness of our KD-QAT methodology on an established chat application, we experiment with the popular LLaMAv2-Chat model (Touvron et al. (2023)), and quantize it to W4A16222WxAy quantization indicates quantizing all the weights and output activations to x and y bits respectively.. For knowledge distillation based QAT, we use full-precision chat model as the teacher and quantized model as our student. Contrary to LLM-QAT (Liu et al. (2023)), we observe that using a weighted sum of cross-entropy and KL divergence for training leads to more stable and better convergence.